OpenStudy (anonymous):

A bag contains 8 red marbles, 4 white marbles, and 5 blue marbles. You draw one marble at random. Find P(red and blue).

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (raden):

draw one marble ?

OpenStudy (anonymous):

?

OpenStudy (raden):

draw one marble, impossible get 2 events

OpenStudy (anonymous):

okay theres 17 marbles total. Am i supposed to take away a marble from the red and blue ?

OpenStudy (raden):

i think ur question is P(red or blue) not P(red and blue)

OpenStudy (anonymous):

You just said the same thing.

OpenStudy (raden):

no, it is different

OpenStudy (anonymous):

A) 40/17 B) 0 C) 12/17 D) 13/16

OpenStudy (raden):

if u draw 2 marbles, yes correct it is P(red and blue) but the condition of this case just only draw one marble, so it P(R or B)

OpenStudy (raden):

P(R or B) = P(R) + P(B) = 8/17 + 5/17 = 12/17

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

A spinner is numbered from 1 through 10 with each number equally likely to occur. What is the probability of obtaining a number less than 2 or greater than 7 in a single spin?

OpenStudy (anonymous):

A) 2/5 B) 1/2 C) 2 2/2 D) 3/10

OpenStudy (raden):

less than 2, it is 1 more than 7, they are 8,9 and 10 so, the total event is 4 events 4 of 10, means 4/10=2/5

OpenStudy (anonymous):

Joey’s sock drawer is unorganized and contains 3 black dress socks, 5 black ankle socks, 7 brown dress socks, and 6 brown ankle socks. What is the probability that Joey will blindly reach into his sock drawer and pull out a sock that is brown or a dress sock?

OpenStudy (anonymous):

A) 23/21 B) 1/3 C) 16/21 D) 4

OpenStudy (raden):

sorry, i have to go now... i want going to pray to mosque :) next time i will help u

OpenStudy (anonymous):

man okay.

OpenStudy (raden):

well, for Q3) P(brown or dress sock) = 13/21 + 10/21 - 7/21 = 16/21 btw for number 1 question before, just want confirm is ur question : P(red and blue) OR P(red or blue) ?? if P(red and blue) then the answer is 0 if P(red or blue) then the answer is 12/17